OpenStudy (anonymous):

2/3 x=18, what is 2(x)-5?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(2/3)x= 18 x= 18(3/2)= 27 2x-5= 2(27)-5= 54-5= 49